Q.1.Casez Ltd., manufacturer of suitcases, makes 700 boxes per annum. Selling price of each suitcase is
Rs.1000 and the variable cost is Rs.400 per piece. It incurs fixed expenses of Rs.1,20,000 per year .Only
10% of its sales are made in cash. Every year the firm spends Rs.20, 000 on advertising and this year its
investment in the purchase of raw materials is 20% of its total turnover. Additional information is given
below.(in Rs. ‘000)
i Opening balance of
a Raw materials, stores, etc. 40
b Work-in-process 35
c Finished goods 80
d Debtors 90
e Creditors 120
ii Closing balance of
a. Raw materials, stores, etc. 45
b. Work-in-process 55
c. Finished goods 90
d. Debtors 100
e. Creditors 70
iii. Depreciation 30
iv. Excise duty 30
v. Selling and administration expenses 40
Assuming uniform production and sales throughout the year, calculate the operating cycle for Casez Ltd.
Answer #1

Operating Cycle= [(365/Purchases)*Average Inventories]+[(365/Receivables)*Average Accounts Receivables]

Purchases=20% of total turnover

Cost of raw materials=700000*20%=140000

Sales=700*1000=700000

Cash Sales= 10% which is 70000

Credit Sales=630000 (700000-70000)

Average Inventory= (Beginning Inventory+Ending Inventory)/2

=40000+45000/2=42500

Average Account Receivables=(Beginning Account Receivable+Ending Account Receivable)/2

=90000+100000/2=95000

Operating Cycle=[(365/Purchases)*Average Inventories]+[(365/Receivables)*Average Accounts Receivables]

=[(365/140000)*42500]+[365/630000)*95000]

=110.80+55.039

Operating Cycle =165.839 Days